OpenStudy (anonymous):

Urgent Help......Pleasee The sum of the series \( \bf 45^2-43^2+44^2-42^2+43^2-41^2+42^2-40^2.....\) to 30 terms is:- a)2000 b)1590 c)2143 d)2220

OpenStudy (anonymous):

When i solve it, i got 7020. Am i right?

OpenStudy (anonymous):

@ash2326 @Callisto @Hero

OpenStudy (anonymous):

@ganeshie8 @kropot72 @skullpatrol

OpenStudy (anonymous):

Help me Pleaseeeeee

OpenStudy (jack1):

i got 2220

OpenStudy (anonymous):

how?

OpenStudy (jack1):

kinda lazy non-maths way: excel n1 =45 n30 =29 sum of the positives = 21940 sum of the negatives = -19720 sum of these = 2220

OpenStudy (anonymous):

Well, I think it's of the form a_n= [(46-n)^2 - (44-n)^2] starting at n=1

OpenStudy (anonymous):

whatever number you plug in for n, all other numbers but the first positive and last negative should cancel since this is a telescoping series.

OpenStudy (anonymous):

help me by applying A.P formula

OpenStudy (callisto):

\[45^2 - 43^2 + 44^2 - 42^2 + 43^2 - 41^2 + 42^2 - 40^2+...\]=(45+43)(45-43) + (44+42)(44-42)+(43+41)(43-41)+(42+40)(42-40)+... =2(88+86+84+82+...) The sum there is a A.P. sum

OpenStudy (anonymous):

I get 3540 when n = 30, but since I don't think they expect you to find that, if you plug in n = 15 (will result in 30 terms total) I get 2220 as well.

OpenStudy (anonymous):

like this:

OpenStudy (anonymous):

@Callisto i got 176,172,168,......

OpenStudy (anonymous):

@hartnn can you help me?

OpenStudy (anonymous):

then, a = 176 and d=-4 in this way, i got 3540

OpenStudy (callisto):

Hmm, make good use of factorization. \(45^2−43^2+44^2−42^2+43^2−41^2+42^2−40^2+...\) <- 30 terms Factorize every TWO terms by the formula \(a^2-b^2 = (a+b)(a+b)\) =(45+43)(45-43) + (44+42)(44-42)+(43+41)(43-41)+(42+40)(42-40)+... <- 15 terms = (88(2) + 86(2) + 84(2) + 82(2) + ... Factorize again by taking out the common factor "2" =2(88+86+84+82+...) There are 15 terms in the sum, and the sum is a AS

OpenStudy (callisto):

AS = arithmetic sequence. I suppose you know how to work out the sum for it

hartnn (hartnn):

maybe you took n=30 actually n=15 as callisto explained

OpenStudy (anonymous):

oh! i got now.....thank you guys....

OpenStudy (jack1):

or you could logic it out: 45 and 44 are the only positive non-repeating numbers in the first 30 terms of the series 30 and 29 are the only negative non-repeating numbers in the first 30 terms of the series all the rest are cancelled out so: 45^2 + 44^2 = 3961 -(30^2)+ -(29^2) = -1741 3961-1741 = 2220

OpenStudy (anonymous):

I literally said that. that's the definition of a telescoping series.

OpenStudy (jack1):

u said the first and the last numbers tho dude, it;s actually the first 2 and the last 2...